OpenStudy (anonymous):

solve the system by substitution -4.5x-2y=-12.5 3.25-y=0.75

OpenStudy (anonymous):

There is no x in the latter equation?

OpenStudy (texaschic101):

is the 2nd one supposed to be : 3.25x - y = 0.75 ??

OpenStudy (texaschic101):

if so... 3.25x - 0.75 = y now sub 3.25x - 0.75 in for y in the other equation....then solve for x

OpenStudy (anonymous):

did you solve it?

OpenStudy (anonymous):

i didnt

OpenStudy (anonymous):

do you still need help with it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve for y in your second equation

OpenStudy (anonymous):

using the above and subbing it for y we get \[\frac{ -9 }{2 }x-2(\frac{ 13 }{ 4 }x-\frac{ 3 }{ 4 })=\frac{ -25 }{ 2 }\] which becomes \[\frac{ -9 }{ 2 }x - \frac{ 13 }{ 2 }x +\frac{ 3 }{ 2 } = \frac{ -25 }{ 2 }\] then we just move things around a bit and combine \[-11x = \frac{ -25 }{ 2 } - \frac{ 3 }{ 2 }\] then it's just \[-11x = -14 -> x = 14/11\]

OpenStudy (anonymous):

from there you can just plug in your x value and solve for y

OpenStudy (anonymous):

\[3.25(\frac{ 14 }{ 11 })-y=0.75 \rightarrow y = 3.25*\frac{ 14 }{ 11 }-0.75\]

OpenStudy (anonymous):

still having a problem?